Opened 10 years ago

Closed 7 years ago

#465 closed flight processing (fixed)

GB12/05, flight day 290/2012, Brampton

Reported by: knpa Owned by: tipo
Priority: alpha 4 medium Milestone: 2012 data processing completion
Component: Archiving Keywords:
Cc: Other processors:

Description (last modified by edfi)

Data location: /users/rsg/arsf/arsf_data/2012/flight_data/uk/GB12_05-2012_290_Brampton

Data arrived from ARSF via SATA disk on 19/10/2012

GB12/05 Scientific objective: unknown

GB12/05 Priority: a4m

PI: Ross Hill

Notes: FW LiDAR present

Sensors:

Camera (09/04/13)
Eagle (07/12/12)
Hawk (07/12/12)
LiDAR (07/12/12)
FW LiDAR (23/07/13)

Change History (38)

comment:1 Changed 10 years ago by knpa

  • Description modified (diff)

comment:2 Changed 10 years ago by knpa

  • Description modified (diff)
  • Summary changed from GB12/05, flight day 289/2012, Brampton to GB12/05, flight day 290/2012, Brampton

comment:3 Changed 10 years ago by edfi

finished navigation processing, created sol files incl. camera.

comment:4 Changed 10 years ago by knpa

Beginning LiDAR processing.

knpa teaching edfi.

comment:5 Changed 10 years ago by edfi

LiDAR classified, no vector data available. LiDAR delivery complete

comment:6 Changed 10 years ago by knpa

Hyperspectral processing (knpa showing tipo)

comment:7 Changed 10 years ago by mark1

Lidar delivery check completed - all looks ok.

comment:8 Changed 10 years ago by tipo

Finished hyperspectral processing.

comment:9 Changed 10 years ago by edfi

Beginning hyperspectral delivery check

comment:10 Changed 10 years ago by edfi

ReadMe missing, DEM needs renaming to convention, check_apl_cmd not working could be similar to 249 (Luton) in that the hyp_genreadme_config file are not correct, in particular the file names do not match the naming convention.

Level 1 files look ok

Will look again when these have been adjusted

comment:11 Changed 10 years ago by tipo

Fixed the above stuff and reprocessed for all bands

comment:12 Changed 10 years ago by edfi

In flightlines/level1b which files are correct e290081b_n_sct-0.01.bil are correct or not needed? The same in line_information and navigation.

Just a versioning issue I think from multiple runs. Everything else looks fine

comment:13 Changed 10 years ago by tipo

Removed the old files and renamed the newer ones to remove the _sct-0.01 bit

comment:14 Changed 10 years ago by edfi

Hyperspectral check complete all now looks fine

comment:15 Changed 10 years ago by edfi

Beginning Camera delivery check

comment:16 Changed 10 years ago by edfi

LiDAR and Hyperspectral loaded onto HDD 142. Camera may be sent at a later date

comment:17 Changed 10 years ago by edfi

LiDAR and Hyperspectral sent special delivery to PI (Ross Hill) 07/12/12

comment:18 Changed 10 years ago by edfi

  • Description modified (diff)

comment:19 Changed 10 years ago by tipo

SCT values:

LineEagleHawk
1-0.03-0.06
2-0.02-0.07
3-0.03-0.06
4-0.01-0.06
5-0.01-0.06
6-0.02-0.06
7-0.01-0.06
8-0.01-0.06

comment:20 Changed 10 years ago by edfi

Starting Vector check for LiDAR and Hyperspectral

comment:21 Changed 10 years ago by edfi

Vector check complete. Total coverage of the area was not possible due to gaps in the data sent by OS. Readme updated. Ready for re-check

comment:22 Changed 10 years ago by anhi

LiDAR Vector Delivery Check:

No problems

comment:23 Changed 10 years ago by anhi

Hyperspectral vector check:

  • gdal_translate example command in Readme needs correcting

comment:24 Changed 10 years ago by edfi

Unzipping files to remake Readme

comment:25 Changed 10 years ago by edfi

New ReadMe made and unzipped files removed

comment:26 Changed 10 years ago by anhi

Readme checked no problems found.

comment:27 Changed 9 years ago by edfi

  • Description modified (diff)

comment:28 Changed 9 years ago by edfi

Available RCD dispatched 09/04/13 - RCD suffered error during flight

comment:29 Changed 9 years ago by tipo

  • Component changed from Processing: general to Archiving
  • Owner set to benj

comment:30 Changed 9 years ago by tipo

  • Owner changed from benj to tipo
  • Status changed from new to assigned

Beginning archiving

comment:31 Changed 9 years ago by edfi

  • Description modified (diff)

comment:32 Changed 9 years ago by edfi

FW LiDAR Processed, ready for check

comment:33 Changed 9 years ago by tipo

LiDAR checked, no problems found.

comment:34 Changed 9 years ago by edfi

  • Description modified (diff)

comment:35 Changed 9 years ago by edfi

FW LiDAR Dispatched to PI

comment:36 Changed 7 years ago by tec

Archiving
Started

Ready for upload

Last edited 7 years ago by tec (previous) (diff)

comment:37 Changed 7 years ago by dac

Archiving

Received conformation from Wendy raw and processed data have been archived 27/07/2015

comment:38 Changed 7 years ago by dac

  • Resolution set to fixed
  • Status changed from assigned to closed
Note: See TracTickets for help on using tickets.